# Exploring the Impact of Nuclear Parton Distribution Functions on Open Heavy Flavor Azimuthal Angular Correlations in Pb-Pb Collisions at $\sqrt{s_{NN}}$ = 5.5 TeV

### v1

@@ Abstract
In this paper, we explore the impact of nuclear parton distribution functions (nPDFs) on the azimuthal angular correlation of open heavy flavor hadrons in Pb-Pb collisions at $\sqrt{s_{{}_{\rm{NN}}}}$ = 5.5 TeV, utilizing PYTHIA8 + Angantyr model and various PDFs. The method involves a transverse momentum ($p_{\rm{T}}$) differential assessment of azimuthal angular correlations between prompt $\rm{D^{0}-\bar{D}^{0}}$, $\rm{B-\bar{B}}$, and non-prompt D-mesons. The results show that nPDFs significantly modify the near-side and away-side peaks of the azimuthal angular correlation distribution. A differential $p_{\rm{T}}$ analysis allowed us to find the interplay between various production mechanisms and the significant impact of nPDFs on the dynamics of heavy flavor hadron production. Apart from the contributions originating from leading order (LO) production processes like pair creation in the away side correlation distribution, the identification of a double peak structure in the near side of azimuthal angular correlation indicates contributions from next-to-leading order (NLO) processes. The paper also emphasizes the relevance of nPDF in the context of heavy quark thermalization in heavy-ion collision analyses. The findings indicate that nPDFs alter charm quark thermalization at low $p_{\rm{T}}$, with negligible effects on beauty quark thermalization. This comprehensive analysis of azimuthal correlations and nPDF effects enhances our understanding of the initial state conditions and the dynamics of heavy flavor production in heavy-ion collisions.
@@ I Introduction
Experiments at the Large Hadron Collider at CERN, Geneva and the Relativistic Heavy Ion Collider (RHIC) at BNL, USA, explore the properties of quark-gluon plasma (QGP), an exotic state of matter, by colliding ultra-relativistic heavy-ions STAR:2005gfr ; Braun-Munzinger:2015hba ; ALICE:2022wpn ; ALICE:2016fzo . The QGP was discovered at the SPS, in Pb-Pb collisions at the center of mass energy $\sqrt{s_{{}_{\rm{NN}}}}$ = 17.3 GeV E877:1994plr ; NA50:2000brc , inspiring future experiments such as ALICE and STAR to perform precise measurements and understand its properties.
@@
In heavy-ion collisions, measurements of heavy flavor nuclear modification factor ($\mathrm{R_{AA}}$:2017qjw ; ALICE:2015zhm ; ALICE:2019hno ; ALICE:2015vxz ; ALICE:2018lyv ; STAR:2014wif and elliptic flow ($\mathrm{v_{2}}$) of prompt and non-prompt charm hadrons ALICE:2023gjj ; STAR:2017kkh ; ALICE:2014qvj ; ALICE:2013olq ; ALICE:2020hdw have provided ample information about energy loss mechanisms and thermalization of heavy quarks in the QGP. The measurement of charm $\mathrm{v_{2}}$ as a function of $p_{\rm{T}}$ by ALICE ALICE:2013olq ; ALICE:2020hdw ; ALICE:2023gjj , quotes their participation in the collective medium, corresponding to its thermalization in the region $82<p_{\mathrm{T}}<$ 8 GeV/$c$. Although this measurement confirms the thermalization of charm quarks in the QGP, it cannot comment on the degree of thermalization viz. fraction of thermalized charm quarks. The azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ can be an excellent observable in this regard, which will allow one to look for effects of thermalization in differential $p_{\rm{T}}$ ranges. The azimuthal angular correlation Beraudo:2014boa ; PHENIX:2018wex ; Zhang:2019bkf ; Zhang:2018ucx ; ALICE:2019oyn ; ALICE:2021kpy ; ALICE:2016clc , is the distribution of the differences in azimuthal angles, $\rm{\Delta\varphi=\varphi_{trig}-\varphi_{assoc}}$, and pseudorapidities, $\rm{\Delta\eta=\eta_{trig}-\eta_{assoc}}$, where $\rm{\varphi_{trig}}$ ($\rm{\eta_{assoc}}$) and $\rm{\varphi_{trig}}$ ($\rm{\eta_{assoc}}$) are the azimuthal angles (pseudorapidities) between the trigger 111 reference particle with respect to which correlation is computed. and associated particles 222  particles being correlated with the trigger., respectively. The correlation function usually contains a ``````` `near-side” (NS) peak and an ``````` `away-side” (AS) peak at $\rm{\Delta\varphi=0}$ and $\rm{\Delta\varphi=\pi}$ respectively over a wide range of $\rm{\Delta\eta}$. For fully thermalized charm quarks, the flight direction will be completely randomized leading to flat $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation ALICE:2022wwr , and any deviation from flatness can be used to estimate the degree of thermalization. Besides thermalization, $\rm{D^{0}-\bar{D}^{0}}$ and $\rm{B-\bar{B}}$ azimuthal angular correlation is an excellent probe of heavy flavor production mechanisms. The LO processes are important for the AS peak, while the NLO process like gluon splitting is a major contributor to the NS of $\rm{D^{0}-\bar{D}^{0}}$ correlation. A $p_{\rm{T}}$ dependent study will allow disentangling the contribution of pair creation and gluon splitting processes in charm production, and comment on the degree of thermalization.
@@
The ALICE3 experiment investigates the thermalization of charm quarks by measuring the azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ ALICE:2022wwr . In this contribution, we study the effects of nPDFs on the azimuthal angular correlation of prompt $\rm{D^{0}-\bar{D}^{0}}$, $\rm{B-\bar{B}}$, and non-prompt $\rm{D-\bar{D}}$ in Pb-Pb collisions at $\sqrt{s_{NN}}=5.5$ TeV. We use the PYTHIA8+Angantyr model Sjostrand:2006za ; Bierlich:2018xfw to study non-collective effects and provide an effective comment on heavy quark thermalization in heavy ion collisions. Because PYTHIA8 lacks a thermal medium, it is an ideal platform to isolate and study the effects of nPDFs, and make conclusive remarks about the thermalization of heavy flavors. Since nPDFs can affect the production of heavy quarks during initial and final state radiations (ISR and FSR) via hard scatterings, their imprints can reflect in the azimuthal angular correlation of open charm and beauty mesons. The comparison of yields obtained from azimuthal angular correlation of $\rm{D^{0}}$ and $\rm{B}$ mesons can give us flavor-dependent and $p_{\rm{T}}$-dependent relative contributions of gluon splitting and pair creation processes. For non-prompt D-mesons, these calculations will serve as a proxy for beauty quarks and allow the investigation of different decay-kinematic effects.
@@
This article is organized as follows, In Section II, we briefly discuss about parton distribution function. Section III contains a detailed discussion of the event generation, methodology, and analysis strategy. Section  IV includes the results of the study. This section is divided into three subsections for prompt $\rm{D^{0}-\bar{D}^{0}}$, $\rm{B-\bar{B}}$, and non-prompt $\rm{D-\bar{D}}$ azimuthal angular correlations respectively. Important findings of the study are summarized in section V.
II Parton distribution function
@@
Parton distribution functions play a pivotal role in hadron and particle physics, as they are crucial for interpreting experimental data from various high-energy processes. PDFs cannot be derived from first principles; instead, they are obtained by comparing theoretical calculations of hadronic cross-sections with experimental data. These functions are modeled using fit parameters optimized through data to achieve matching theoretical predictions. At an input interaction scale $Q^{2}_{0}$ of partons, PDFs are parameterized for each parton via the equation:
@@
As previously mentioned, the parameterization of PDFs is set at an input scale $Q_{0}$; these PDFs subsequently evolve to the scale of data $Q$ using the DGLAP equation Gribov:1972ri ; Lipatov:1974qm ; Altarelli:1977zs ; Dokshitzer:1977sg . Predictions for hadronic cross-sections are obtained by convoluting the evolved parton distribution functions with partonic cross-sections at a given perturbative order. Optimal PDF parameters are found using a suitable optimization method, commonly the log-likelihood or $\chi^{2}$ minimization technique.
@@
$f^{(A,Z)}(x,Q^{2})=\frac{Z}{A}f^{p/A}(x,Q^{2})+\frac{(A-Z)}{A}f^{n/A}(x,Q^{2})$ (3)
In this expression, $\it{A}$ is the atomic mass number, $\it{Z}$ is the atomic number, and $f^{p,n/A}$ are the proton and neutron-bound PDFs, related to the nucleon PDFs by
$f^{p,n/A}(x,Q^{2})=R^{A}_{f}(x,Q^{2})f^{p,n}(x,Q^{2})$ (4)
where $R^{A}_{f}(x,Q^{2})$ is a scale-dependent nuclear modification. This factor illuminates the various nuclear modifications concerning the free proton in different $x$ regions. A shadowing behavior ($R<1$) in the low $x\leq 0.1$ area has been observed in experimental data. This is followed by anti-shadowing ($R>1$) around $x\approx 0.1$ and the EMC effect ($R<1$) in the valence region, $x\approx 0.4$ EuropeanMuon:1992pyr . The mechanisms generating these effects are still under investigation but appear as a general feature in all nuclear PDFs.
The free proton PDFs under our consideration are NNPDF2.3 NNPDF:2017mvq and CT18Anlo Hou:2019efy PDF. The NNPDF2.3 is a default proton PDF for which different PYTHIA8 processes are tuned, providing a standard PYTHIA8 baseline for comparison. It is the first global PDF set to systematically include all relevant LHC data up to 2012, employing a variable flavor number scheme based on FONLL  Cacciari:1998it for heavy quarks. The CT18 QCD global fit uses high-statistics LHC datasets, HERA I + II DIS data, and datasets from the CT14 global QCD analysis Dulat:2015mca , totaling 40 datasets. The global fit also incorporates ATLAS $\sqrt{s}=7$ TeV $W$ and $Z$ boson data. For CT18Anlo, heavy quarks are parameterized using the SACOT-$\chi$ scheme Aivazis:1993pi .
The nuclear PDFs used are EPPS21 Eskola:2021nhw and nNNPDF3 AbdulKhalek:2022fyi . The EPPS21 nuclear PDF is derived using CT18nlo as the free proton PDF, with nuclear modifications parameterized at the charm pole mass threshold $Q_{0}=m_{\mathrm{charm}}=1.3$ GeV. This parameterization addresses the EMC effect, anti-shadowing, and shadowing at appropriate $x$ values. The number of free parameters in this system is six, and heavy quarks are treated using the SACOT-$\chi$ scheme. The nNNPDF3 uses Machine Learning to extract PDFs from the data, based on NNPDF4 NNPDF:2023tyk as the free proton PDF, with heavy quarks treated using the FONLL scheme. The use of nNNPDF3.0 is important for examining the effects of different heavy quark treatments.
@@ III.1 The Angantyr model
The PYTHIA8 event generator Sjostrand:2006za ; Singh:2021edu ; Sjostrand:2007gs ; Bierlich:2018xfw ; Skands:2014pea ; Campbell:2022qmc ; Corke:2010yf ; Lonnblad:2021fyl is used to investigate pp, ep, p-A, and A-A collisions in depth. It uses 2 → 2 QCD matrix elements calculated with leading-order accuracy, incorporating next-to-leading order effects during the parton showering stage. Leading-logarithmic $p_{\rm{T}}$ ordering governs the parton showering, and an extra veto prevents soft-gluon emission divergences. The hadronization process is managed using the Lund string-fragmentation model. It offers a plethora of processes and tunes to use based on the physics involved in the study. PYTHIA8 employs an initial hard scattering with the incoming proton beams, followed by ISR, multi-parton interactions (MPI) Corke:2009tk ; Diehl:2011yj ; Sjostrand:2017cdm and FSR processes, comprising the underlying event. The high $p_{\rm{T}}$ partons create showers or jets that fragment and hadronize based on the Lund string fragmentation model Andersson:1983ia . Hadronization is accomplished through the Color Reconnection (CR) mechanism between fragmenting partons Khoze:1994fu ; Gieseke:2012ft ; Rathsman:1998tp ; Bierlich:2015rha , which is accomplished by rearranging the strings between them. This affects hadronization by modifying the total string length. When the string length is small enough after the subsequent creation of quark-antiquark combinations, the partons hadronize into hadrons. The MPI and CR phenomena in PYTHIA8 play an essential role in the particle production, as evidenced by the charged-particle multiplicity and $p_{\rm{T}}$ distributions, and particle ratios for pp collisions at LHC energies Christiansen:2015yqa .
@@
In PYTHIA8, heavy-flavors are produced via leading order perturbative scattering processes such as gluon fusion ($gg\rightarrow Q\overline{Q}$) or pair annihilation ($q\overline{q}\rightarrow Q\overline{Q}$). PYTHIA8 approximates higher-order contributions inside its LO framework via flavor excitations ($gQ\rightarrow Qg$) or gluon splittings ($g\rightarrow Q\overline{Q}$) that give rise to heavy-flavor generation during high $p_{\rm{T}}$ parton showers Braun-Munzinger:2003pwq ; Norrbin:2000zc .
@@
In PYTHIA8, heavy-ion collisions are simulated using the Angantyr model Bierlich:2018xfw , which extrapolates proton-proton (pp) collision dynamics to nucleus-nucleus (AA) collisions. This generalization is achieved as follows. The positions of nucleons within the nucleus are determined by generating a nuclear density distribution using Monte Carlo methods. The Glauber formalism, which employs the eikonal approximation in impact parameter space, is used to calculate the number of interacting nucleons and binary collisions. In this case, projectile nucleons are supposed to travel in straight lines, colliding with nucleons in the target several times. The contribution from individual nucleon-nucleon interactions in the final state is estimated using the Fritiof model, which accounts for both diffractively and non-diffractively wounded nucleons and manages soft particle production Bierlich:2016smv . Hard particles are produced in individual nucleon-nucleon sub-collisions. At high energies, hard partonic sub-collisions play a crucial role, which is addressed by introducing the concepts of primary and secondary absorptive interactions. For both absorptive and diffractive interactions, PYTHIA8 generates the relevant nucleon-nucleon parton-level events utilizing the entire multi-parton interaction (MPI) machinery.
@@
While the Angantyr model does not employ a thermalized medium, it accurately describes the multiplicity of heavy-ion collisions as a function of centrality and y distribution. As a result, Angantyr offers the unique opportunity to categorize the collective and non-collective effects in heavy-ion collisions. As $\rm{D^{0}-\bar{D}^{0}}$ correlation measurements are sensitive to heavy quark thermalization and its participation in the QGP for $p_{\rm{T}}$ ranges from $\mathrm{2-6}$ GeV/$c$, our study will allow separating the non-collective components of heavy-flavor production and dynamics in heavy-ion collisions.

### v2

# An Impact of Parton Distribution Functions on Open Heavy Flavor Azimuthal Angular Correlations in Pb-Pb Collisions at $\sqrt{s_{NN}}$ = 5.5 TeV

@@ Abstract
In this paper, we investigate the impact of modern free parton distribution functions and nuclear parton distribution functions on the azimuthal angular correlation of open heavy flavor hadrons in Pb-Pb collisions at $\sqrt{s_{{}_{\rm{NN}}}}$ = 5.5 TeV using PYTHIA8 + Angantyr model. The method involves a transverse momentum ($p_{\rm{T}}$) differential assessment of azimuthal angular correlations between prompt $\rm{D^{0}-\bar{D}^{0}}$ and $\rm{B-\bar{B}}$. By analyzing the $p_{\rm{T}}$-dependent behavior of these correlations, we probe the interplay between different production mechanisms and assess the impact of parton distribution functions on heavy-flavor hadron production dynamics. Our results indicate that nuclear parton distribution functions significantly modify both near-side and away-side peaks of the azimuthal correlation distribution compared to the default PYTHIA8 baseline, particularly at low $p_{\rm{T}}$. The near-side double-peak structure suggests contributions from next-to-leading order (NLO) processes, such as gluon splitting, and its presence is strongly correlated with the number of hard multi-partonic interactions. Additionally, by quantitatively comparing the change in correlation widths across various PDF settings, we comment on the role of nuclear parton distribution functions in heavy-quark thermalization in heavy-ion collisions. This comprehensive study of azimuthal correlations enhances our understanding of initial-state conditions and heavy-flavor production dynamics in nuclear collisions.
@@ I Introduction
Experiments at the Large Hadron Collider at CERN, Geneva, and the the Relativistic Heavy Ion Collider (RHIC) at BNL, USA, explore the properties of quark-gluon plasma (QGP), an exotic state of matter, by colliding ultra-relativistic heavy-ions STAR:2005gfr ; Braun-Munzinger:2015hba ; ALICE:2022wpn ; ALICE:2016fzo . The QGP was discovered at the SPS, in Pb-Pb collisions at center of mass energy $\sqrt{s_{{}_{\rm{NN}}}}$ = 17.3 GeV NA50:2000brc , inspiring future experiments such as ALICE and STAR to perform precise measurements and understand its properties.
@@
In heavy-ion collisions, measurements of heavy flavor nuclear modification factor ($\mathrm{R_{AA}}$:2017qjw ; ALICE:2015zhm ; ALICE:2019hno ; ALICE:2015vxz ; ALICE:2018lyv ; STAR:2014wif and elliptic flow ($\mathrm{v_{2}}$) of prompt and non-prompt charm hadrons ALICE:2023gjj ; STAR:2017kkh ; ALICE:2014qvj ; ALICE:2013olq ; ALICE:2020hdw have provided ample information about energy loss mechanisms and thermalization of heavy quarks in the QGP. The measurement of charm $\mathrm{v_{2}}$ as a function of $p_{\rm{T}}$ by ALICE ALICE:2013olq ; ALICE:2020hdw ; ALICE:2023gjj , quotes their participation in the collective medium, corresponding to its thermalization in the region $82<p_{\mathrm{T}}<$ 8 GeV/$c$. Although this measurement confirms the thermalization of charm quarks in the QGP, it cannot comment on the degree of thermalization, viz. fraction of thermalized charm quarks. The ALICE3 experiment investigates the thermalization of charm quarks by measuring the azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ ALICE:2022wwr . The azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$ can be an excellent observable in this regard, which will allow one to look for effects of thermalization in differential $p_{\rm{T}}$ ranges. The azimuthal angular correlation Beraudo:2014boa ; PHENIX:2018wex ; Zhang:2019bkf ; Zhang:2018ucx ; ALICE:2019oyn ; ALICE:2021kpy ; ALICE:2016clc , is the distribution of the differences in azimuthal angles, $\rm{\Delta\varphi=\varphi_{trig}-\varphi_{assoc}}$, and pseudorapidities, $\rm{\Delta\eta=\eta_{trig}-\eta_{assoc}}$, where $\rm{\varphi_{trig}}$ ($\rm{\eta_{trig}}$) and $\rm{\varphi_{assoc}}$ ($\rm{\eta_{assoc}}$) are the azimuthal angles (pseudorapidities) of the trigger 111 reference particle with respect to which correlation is computed. and associated particles 222  particles being correlated with the trigger., respectively. The structure of the correlation function usually contains a ``````` `near-side” (NS) peak and an ``````` `away-side” (AS) peak at $\rm{\Delta\varphi=0}$ and $\rm{\Delta\varphi=\pi}$ respectively over a wide range of $\rm{\Delta\eta}$. For fully thermalized charm quarks, the flight direction will be completely randomized, leading to flat $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation ALICE:2022wwr , and any deviation from flatness can be used to estimate the degree of thermalization. Besides thermalization, $\rm{D^{0}-\bar{D}^{0}}$ and $\rm{B-\bar{B}}$ azimuthal angular correlation is an excellent probe of heavy flavor production mechanisms. The LO processes are important for the AS peak, while the NLO process like gluon splitting is a major contributor to the NS of $\rm{D^{0}-\bar{D}^{0}}$ correlation. A $p_{\rm{T}}$ dependent study will allow disentangling the contribution of pair creation and gluon splitting processes in charm production and comment on the degree of thermalization.
@@
In this contribution, we study the effects of nPDFs on the azimuthal angular correlation of prompt $\rm{D^{0}-\bar{D}^{0}}$, and $\rm{B-\bar{B}}$ in Pb-Pb collisions at $\sqrt{s_{NN}}=5.5$ TeV. We use the PYTHIA8+Angantyr model Sjostrand:2006za ; Bierlich:2018xfw to study non-collective effects and provide an effective comment on heavy quark thermalization in heavy ion collisions. Because PYTHIA8 lacks a thermal medium, it is an ideal platform to isolate and study the effects of nPDFs, and make conclusive remarks about the thermalization of heavy flavors. Since nPDFs can affect the production of heavy quarks during initial and final state radiations (ISR and FSR) via hard scatterings, their imprints can reflect in the azimuthal angular correlation of open charm and beauty mesons. The comparison of yields obtained from azimuthal angular correlation of $\rm{D^{0}}$ and $\rm{B}$ mesons can give us flavor-dependent and $p_{\rm{T}}$-dependent relative contributions of gluon splitting and pair creation processes.
@@
This article is organized as follows, In Section II, we briefly discuss about PDFs. Section III contains a detailed discussion of the event generation, methodology, and analysis strategy. Section  IV includes the results of the study. Important findings of the study are summarized in section V.
II Parton distribution functions
@@
Parton distribution functions play a pivotal role in hadron and particle physics, as they are crucial for interpreting experimental data from various high-energy processes. PDFs cannot be derived from first principles; instead, they are determined by comparing theoretical predictions of hadronic cross-sections with experimental data. These functions are modeled using fit parameters optimized through data to achieve matching theoretical predictions. At an input interaction scale $Q^{2}_{0}$ of partons, PDFs are parameterized for each parton via the equation:
@@
As previously mentioned, the parameterization of PDFs is set at an input scale $Q_{0}$; these PDFs subsequently evolve to the scale $Q$ using the DGLAP equation Gribov:1972ri ; Lipatov:1974qm ; Altarelli:1977zs ; Dokshitzer:1977sg . Predictions for hadronic cross-sections are obtained by convoluting the evolved PDFs with partonic cross-sections at a given perturbative order. Optimal PDF parameters are found using a suitable optimization method, commonly the log-likelihood or $\chi^{2}$ minimization technique.
@@
$f^{(A,Z)}_{i}(x,Q^{2})=\frac{Z}{A}f^{p/A}_{i}(x,Q^{2})+\frac{(A-Z)}{A}f^{n/A}_{i}(x,Q^{2})$ (3)
In this expression, the pair of atomic numbers $(A,Z)$ identifies the nuclear isotopes, and $f^{p,n/A}_{i}$ are the proton and neutron-bound PDFs, related to the nucleon PDFs by
$f^{p,n/A}_{i}(x,Q^{2})=R^{A}_{i}(x,Q^{2})f^{p,n}_{i}(x,Q^{2})$ (4)
where $R^{A}_{i}(x,Q^{2})$ is a scale-dependent nuclear modification. This factor illuminates the various nuclear modifications concerning the free proton in different $x$ regions. Experimental data have indicated a shadowing behavior ($R<1$) in the low $x<0.1$ region, followed by anti-shadowing ($R>1$) around $x\approx 0.1$, and the EMC effect ($R<1$) in the valence region, $x\approx 0.4$ EuropeanMuon:1992pyr . The mechanisms generating these effects are still under investigation but appear as a general feature in all nuclear PDFs.
The free proton PDFs under our consideration are NNPDF2.3 Ball:2012cx , NNPDF4.0 NNPDF:2021njg and CT18ANLO Hou:2019efy PDF. The NNPDF2.3 is a default proton PDF for which different PYTHIA8 processes are tuned, providing a standard PYTHIA8 baseline for comparison. It is the first global PDF set to systematically include all relevant LHC data up to 2012, employing a variable flavor number scheme based on FONLL  Cacciari:1998it for heavy quarks. The CT18 QCD global fit uses high-statistics LHC datasets, HERA I + II DIS data, and datasets from the CT14 global QCD analysis Dulat:2015mca , in total 40 datasets. The global fit also incorporates ATLAS at $\sqrt{s}=7$ TeV $W$ and $Z$ boson data. For CT18ANLO, parton PDFs are parameterized using the SACOT-$\chi$ scheme Aivazis:1993pi .
The nuclear PDFs used are EPPS21 Eskola:2021nhw and nNNPDF3 AbdulKhalek:2022fyi . The EPPS21 nuclear PDF is derived using CT18ANLO as the free proton PDF, with nuclear modifications parameterized at the charm pole mass threshold $Q_{0}=m_{\mathrm{charm}}=1.3$ GeV. This parameterization addresses the EMC effect, anti-shadowing, and shadowing at appropriate $x$ values. The number of independent PDFs are six, and heavy quarks are treated using the SACOT-$m_{T}$ general-mass variable flavor number scheme (GM-VFNS). The nNNPDF3 uses Machine Learning to extract PDFs from the data, based on NNPDF4 as the free proton PDF, with heavy quarks treated using the FONLL scheme. The use of nNNPDF3.0 is important for examining the effects of different heavy quark treatments. Both nuclear PDFs use a lot of new data to constrain gluon PDFs compared to their predecessors. The most relevant data related to heavy quark production is the measurement of the nuclear modification factor of $\rm{D^{0}}$ mesons in p-Pb collisions by LHCb LHCb:2017yua ; LHCb:2022dmh . Since the production of charm quark is a sensitive probe of gluon PDFs, it allows to put a stringent constraint on the gluon PDFs up to $\rm{x\approx 10^{-5}}$.
@@ III.1 The Angantyr model
The PYTHIA8 event generator Sjostrand:2006za ; Singh:2021edu ; Sjostrand:2007gs ; Bierlich:2018xfw ; Skands:2014pea ; Campbell:2022qmc ; Corke:2010yf ; Lonnblad:2021fyl is used to investigate pp, ep, p-A, and A-A collisions in depth. It uses 2 → 2 QCD matrix elements calculated with leading-order accuracy, incorporating next-to-leading order effects during the parton showering stage. The parton showering follows a leading-logarithmic $p_{\rm{T}}$ ordering, with soft-gluon emission divergences prevented by an additional veto. The hadronization process is managed using the Lund string-fragmentation model. It offers a plethora of processes and tunes to use based on the physics involved in the study. PYTHIA8 employs an intial hard scattering with the incoming proton beams, followed by ISR, multi-parton interactions (MPI) Corke:2009tk ; Diehl:2011yj ; Sjostrand:2017cdm and FSR processes, comprising the underlying event. The high $p_{\rm{T}}$ partons give rise to showers or jets that fragment and hadronize according to the Lund string fragmentation model Andersson:1983ia . Hadronization is accomplished through the Color Reconnection (CR) mechanism between fragmenting partons Khoze:1994fu ; Gieseke:2012ft ; Rathsman:1998tp ; Bierlich:2015rha , which is accomplished by rearranging the strings between them. This modifies the total string length, affecting hadronization. When the string length is small enough after the subsequent creation of quark-antiquark combinations, the partons hadronize into hadrons. The MPI and CR phenomena in PYTHIA8 play an essential role in the particle production, as evidenced by the charged-particle multiplicity and $p_{\rm{T}}$ distributions, and particle ratios for pp collisions at LHC energies Christiansen:2015yqa .
@@
Leading order perturbative scattering processes of gluon fusion ($gg\rightarrow Q\overline{Q}$) or pair annihilation ($q\overline{q}\rightarrow Q\overline{Q}$) is used for the production of heavy-flavors in PYTHIA8. PYTHIA8 also approximates certain higher-order contributions within its LO framework via flavor excitations ($gQ\rightarrow Qg$), or gluon splittings ($g\rightarrow Q\overline{Q}$) which give rise to heavy-flavor production during high $p_{\rm{T}}$ parton showers Braun-Munzinger:2003pwq ; Norrbin:2000zc .
@@
In PYTHIA8, heavy-ion collisions are simulated using the Angantyr model Bierlich:2018xfw , which extrapolates proton-proton (pp) collision dynamics to nucleus-nucleus (AA) collisions. The positions of nucleons within the nucleus are determined by generating a nuclear density distribution using Monte Carlo methods. The number of interacting nucleons and binary collisions is calculated using the Glauber formalism, which is based on the eikonal approximation in impact parameter space. Here, the projectile nucleons are assumed to travel along straight lines, undergoing multiple sub-collisions with nucleons in the target. The contribution from individual nucleon-nucleon interactions in the final state is estimated using the Fritiof model, which accounts for both diffractively and non-diffractively wounded nucleons and manages soft particle production Bierlich:2016smv . Hard particles are produced in individual nucleon-nucleon sub-collisions. At high energies, hard partonic sub-collisions play a crucial role, which is addressed by introducing the concepts of primary and secondary absorptive interactions. The corresponding nucleon-nucleon parton-level events are generated using the full multi-parton interaction (MPI) machinery in PYTHIA8 for both absorptive and diffractive interactions.
@@
While the Angantyr model does not employ a thermalized medium, it accurately describes the multiplicity of heavy-ion collisions as a function of centrality and rapidity (y) distribution. As a result, Angantyr offers the unique opportunity to categorize the collective and non-collective effects in heavy-ion collisions. As $\rm{D^{0}-\bar{D}^{0}}$ correlation measurements are sensitive to heavy quark thermalization in the QGP, our study will allow separating the non-collective components of heavy-flavor production and dynamics in heavy-ion collisions.
Figure 1: The ratio of $Z$-boson yield (black points) as a function of rapidity measured by ALICE and ATLAS in Pb-Pb collisions at $\sqrt{s}=5.02$ TeV to free-nucleon PDF (CT14). The ratios are compared to predictions using different free proton PDFs (CT14NLO, CT18ANLO) and nuclear PDFs (EPPS16, EPPS21).
Figure 2: Azimuthal angular correlations of prompt $\rm{D^{0}-\bar{D}^{0}}$ mesons from the same hard scattering process in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV, simulated using PYTHIA8+Angantyr with different PDF and nPDF settings. Results are shown for various $p_{\rm{T}}^{\rm D^{0}}$ and $p_{\rm{T}}^{\rm\bar{D^{0}}}$ ranges.